# In Situ Training of Implicit Neural Compressors for Scientific Simulations via Sketch-Based Regularization

### v1

@@ 2.2 JL Regularization for In Situ Learning
We refer to [46] for an extensive treatment of manifolds, but it is sufficient for our purposes to understand that a Riemannian manifold can be viewed as a subset of Euclidean space and behaves locally like a vector space. Informally, we are considering (possibly) lower-dimensional geometric structures embedded in some ambient subset of $\mathbb{R}^{N}$.
Theorem 2 (Manifold JL Transform [47]).
@@
The volume, condition number, and geodesic covering regularity (the latter two of which are defined in [47]) together define the topological regularity of the manifold. Note that this result, as originally stated, includes a factor of $\sqrt{k/N}$ as the sketch is not assumed to be normalized. Theorem˜2 presents a very similar result to Theorem˜1, but where the size of the sketch depends on geometric properties of the manifold, most importantly the intrinsic dimension, and not on the collection of points being sketched. This means that for low-dimensional manifolds of a potentially high-dimensional ambient space, one can construct a small JL transform for an arbitrarily large number of points. The size $k$ depends only logarithmically on the ambient dimension $N$, and in fact, this dependence can be removed completely as shown in subsequent works [48]. This ability to construct JL transformations for manifolds is key to our main result, which we present next.
@@
The proof of the second inequality is motivated by the informal proof of Theorem 2.14 in [49]. Also, note that we construct a JL transform for the shifted manifold $\mathcal{M}-\{\bm{U}_{t}\}$, which is just a convenience to simplify the proof, but does not introduce any complexity to the sketch itself. The shifted manifold will have the same geometric properties as $\mathcal{M}$ that impact the construction of the sketch.
@@ 3 Experiments
We conduct our experiments on the datasets listed in Table˜1, which highlight several potentially challenging or important attributes. The Ignition dataset describes a fully resolved gas wave front on a 2D uniform grid that transitions from a transport phase to a steady state jet. The Neuron dataset describes a 3D diffusion process in an unstructured, non-Cartesian branching neuron tree [50]. The Channel dataset, from the Johns Hopkins Turbulence Database [51, 52], describes 3D turbulent flow on a non-uniform grid. It has been trimmed in space to a $64^{3}$ volume and in time to the first 500 snapshots. We define the notion of a “sample factor” as the relative size of the sketch to the mesh size for a given dataset, i.e., $k/n$.
@@
Our work is implemented in PyTorch [53] and available as an open-source GitHub repository [54]. All training is conducted on a single GPU (A100 or L40S) using the RAdam optimizer, a learning rate of $10^{-4}$, and in single precision (i.e., 32-bit floating point) [55] for both the parameters and data. Given that our goal is to compress the given data, we do not seek to generalize to unseen data; thus, overfitting is not an issue. Our method is fully in situ, so we perform no pre-processing of the data. For all in situ results, the size of the full buffer is always $T_{f}=1$ (only one full snapshot available at a time), and the size of the sketch buffer is $T_{s}=T-1$ (one less than the full time horizon). We do not require sketching the final snapshot because the training loop ends after the last full snapshot has been encoded. The batch size of the sketch buffer and the number of epochs per full snapshot vary between datasets, but are in the range 25–45 and 300–500, respectively.
@@ 3.1 Model Architecture
Within the sine-layer blocks, we use sinusoidal activation functions with the initialization developed in [3]. We also employ skip connections similar to those of [5, 7]. Thus, our networks are defined by two hyperparameters: the depth, i.e., the number of blocks, and the width of the internal linear layers. Because we are targeting the in situ setting, we cannot effectively normalize the data to ensure it is in some specific range, so no activation is used at the final linear layer. This ensures that the target network can produce any signal value. The lack of activation is also useful for the hypernetwork, as we construct the initial output of the hypernetwork to match the initialization of the target INR. To accomplish this, we use the strategy of [6], wherein the bias of the final linear layer in the hypernetwork is set to the exact values given by the initialization of the target INR, and the weights of the last linear layer are scaled by a relatively small factor.
@@ 3.2 Results
Table 2: Performance metrics on test datasets Ignition, Neuron, and Channel at 5%, 1%, and 2% sketch sample factors ($k/n$ where relevant), respectively. Metric statistics are reported as mean plus-or-minus standard deviation from five identical trials. The numbers in the parentheses of the first row of each table are compression ratios.
Table˜2 presents PSNR and RFE performance results on all three datasets in the offline and in situ settings. The key observation to be made is that sketching-based regularization often yields offline levels of performance in the in situ regime. In particular, sketching with the FJLT can approximately match the offline results across all datasets. While subsample sketching performs equally well on the Channel and almost as good on the Ignition datasets, its performance on the Neuron dataset is striking. Although it is not obvious from the presented statistics, the subsample results on the Neuron dataset are partially skewed by a few catastrophic failures. However, even in successful runs, the results are not at FJLT levels. Overall, we believe this result is not too surprising, as the non-Cartesian geometry of the neuron tree makes subsampling a particularly poor choice for sketching.
@@
In Table˜3, we explore how well the theory from Section˜2.2 aligns with our empirical results. Although unreported in Table˜2, we use Offline-FJLT and Offline-Baseline results for Eq.˜6 to compute the $\frac{1+\epsilon}{1-\epsilon}$ ratio from Theorem˜3, which we subsequently use to solve for $\epsilon$. We also compute principal component analysis (lPCA [56]) based estimates of the local manifold dimension for the INC networks used with each of the target datasets. This local dimensionality estimate yields a rough idea of the overall manifold dimension. We perform this analysis using the scikit-dimension package [57] by taking perturbations of size $10^{-5}$ around a nominal point of network parameters. This nominal state is obtained within the in situ setting after training on the first snapshot, i.e., the first time sketching is used.
Datasets Full Loss Sketch Loss Est. Manifold Dim. $M$ Est. Sample Factor $k/n$
@@
Figure 5: Reconstruction comparison of the Ignition data at snapshot $t=13$. The absolute error is computed with respect to the original image, which is presented for reference. The colorbar only references the error cutouts.
@@
Per-snapshot relative error (Eq.˜6) curves for training and testing on the Ignition dataset are given in Fig.˜7. The training loss is split according to Eq.˜4 into the two components $\mathcal{L}_{full}$ and $\mathcal{L}_{sketch}$ corresponding to the two buffers. Thus, the "Full Loss" can be viewed as the instantaneous loss at the current snapshot, while the "Sketch Loss" is the averaged loss over previous sketched snapshots. The "Test Loss" is simply the instantaneous loss for a fixed snapshot after the network has been trained. For early snapshots, before approximately $t=50$, one can observe that the testing error is significantly better than the training loss, indicating the sketched snapshots have also actively been used for learning instead of only preventing forgetting.

### v2

@@ 2.2 JL Regularization for In Situ Learning
We refer to [Lee_manifoldBook2ndEd] for an extensive treatment of manifolds, but it is sufficient for our purposes to understand that a Riemannian manifold can be viewed as a subset of Euclidean space and behaves locally like a vector space. Informally, we are considering (possibly) lower-dimensional geometric structures embedded in some ambient subset of $\mathbb{R}^{N}$.
Theorem 2 (Manifold JL Transform [baraniuk2009random]).
@@
The volume, condition number, and geodesic covering regularity (the latter two of which are defined in [baraniuk2009random]) together define the topological regularity of the manifold. Note that this result, as originally stated, includes a factor of $\sqrt{k/N}$ as the sketch is not assumed to be normalized. Theorem˜2 presents a very similar result to Theorem˜1, but where the size of the sketch depends on geometric properties of the manifold, most importantly the intrinsic dimension, and not on the collection of points being sketched. This means that for low-dimensional manifolds of a potentially high-dimensional ambient space, one can construct a small JL transform for an arbitrarily large number of points. The size $k$ depends only logarithmically on the ambient dimension $N$, and in fact, this dependence can be removed completely as shown in subsequent works [ArminMike2015]. This ability to construct JL transformations for manifolds is key to our main result, which we present next.
@@
The proof of the second inequality is motivated by the informal proof of Theorem 2.14 in [woodruff2014sketching]. Also, note that we construct a JL transform for the shifted manifold $\mathcal{M}-\{\bm{U}_{t}\}$, which is just a convenience to simplify the proof, but does not introduce any complexity to the sketch itself. The shifted manifold will have the same geometric properties as $\mathcal{M}$ that impact the construction of the sketch.
@@ 3 Experiments
We conduct our experiments on the datasets listed in Table˜1, which highlight several potentially challenging or important attributes. The Ignition dataset describes a fully resolved gas wave front on a 2D uniform grid that transitions from a transport phase to a steady state jet. The Neuron dataset describes a 3D diffusion process in an unstructured, non-Cartesian branching neuron tree [Angran3D]. The Channel dataset, from the Johns Hopkins Turbulence Database [channel_flow_1, channel_flow_2], describes 3D turbulent flow on a non-uniform grid. It has been trimmed in space to a $64^{3}$ volume and in time to the first 500 snapshots. We define the notion of a “sample factor” as the relative size of the sketch to the mesh size for a given dataset, i.e., $100\cdot k/n$.
@@
Our work is implemented in PyTorch [pytorch] and available as an open-source GitHub repository [paper-repo]. All training is conducted on a single GPU (A100 or L40S) using the RAdam optimizer, a learning rate of $10^{-4}$, and in single precision (i.e., 32-bit floating point) [single_precision] for both the parameters and data. Given that our goal is to compress the given data, we do not seek to generalize to unseen data; thus, overfitting is not an issue. Our method is fully in situ, so we perform no pre-processing of the data. For all in situ results, the size of the full buffer is always $T_{f}=1$ (only one full snapshot available at a time), and the size of the sketch buffer is $T_{s}=T-1$ (one less than the full time horizon). We do not require sketching the final snapshot because the training loop ends after the last full snapshot has been encoded. The batch size of the sketch buffer and the number of epochs per full snapshot vary between datasets, but are in the range 25–45 and 300–500, respectively.
@@ 3.1 Model Architecture
Within the sine-layer blocks, we use sinusoidal activation functions with the initialization developed in [sitzmann2020implicit]. We also employ skip connections similar to those of [pan2023neural, sales2024implicit]. Thus, our networks are defined by two hyperparameters: the depth, i.e., the number of blocks, and the width of the internal linear layers. Because we are targeting the in situ setting, we cannot effectively normalize the data to ensure it is in some specific range, so no activation is used at the final linear layer. This ensures that the target network can produce any signal value. The lack of activation is also useful for the hypernetwork, as we construct the initial output of the hypernetwork to match the initialization of the target INR. To accomplish this, we use the strategy of [han2023kd], wherein the bias of the final linear layer in the hypernetwork is set to the exact values given by the initialization of the target INR, and the weights of the last linear layer are scaled by a relatively small factor.
@@ 3.2 Results
Table 2: Performance metrics on test datasets Ignition, Neuron, and Channel at 5%, 1%, and 2% sketch sample factors (where relevant), respectively. Metric statistics are reported as mean plus-or-minus standard deviation from five identical trials. The numbers in the parentheses of the first row of each table are compression ratios.
Table˜2 presents PSNR and RFE performance results on all three datasets in the offline and in situ settings. The key observation to be made is that sketching-based regularization often yields offline levels of performance in the in situ regime. In particular, sketching with the FJLT can yield results that approximately match those obtained offline across all datasets. While subsample sketching performs equally well on the Channel and almost as well on the Ignition datasets, its performance on the Neuron dataset is striking. Although it is not obvious from the presented statistics, the subsample results on the Neuron dataset are partially skewed by a few catastrophic failures. However, even in successful runs, the results are not at FJLT levels. Overall, we believe this result is not too surprising, as the non-Cartesian geometry of the neuron tree makes subsampling a particularly poor choice for sketching.
@@
In Table˜3, we explore how well the theory from Section˜2.2 aligns with our empirical results. Although unreported in Table˜2, we use Offline-FJLT and Offline-Baseline results for Eq.˜6 to compute the $\frac{1+\epsilon}{1-\epsilon}$ ratio from Theorem˜3, which we subsequently use to solve for $\epsilon$. We also compute principal component analysis (lPCA [fukunaga1971algorithm]) based estimates of the local manifold dimension for the INC networks used with each of the target datasets. This local dimensionality estimate yields a rough idea of the overall manifold dimension. We perform this analysis using the scikit-dimension package [scikit-dimension] by taking perturbations of size $10^{-5}$ around a nominal point of network parameters. This nominal state is obtained within the in situ setting after training on the first snapshot, i.e., the first time sketching is used.
Datasets Full Loss Sketch Loss Est. Manifold Dim. $M$ Est. Sample Factor (%)
@@
Figure 5: Reconstruction comparison of the Ignition data at snapshot $t=13$. The absolute error in the cutout is computed with respect to the original image, which is presented for reference. The colorbar only references the error cutouts.
@@
Per-snapshot relative error (Eq.˜6) curves for training and testing on the Ignition dataset are given in Fig.˜7. The training loss is split according to Eq.˜4 into the two components $\mathcal{L}_{\text{full}}$ and $\mathcal{L}_{\text{sketch}}$ corresponding to the two buffers. Thus, the “Full Loss” can be viewed as the instantaneous loss at the current snapshot, while the “Sketch Loss” is the averaged loss over previous sketched snapshots. The “Test Loss” is simply the instantaneous loss for a fixed snapshot after the network has been trained. For early snapshots, before approximately $t=50$, one can observe that the testing error is significantly better than the training loss, indicating the sketched snapshots have also actively been used for learning instead of only preventing forgetting.